OpenStudy (anonymous):

A garden table and a bench cost $652 combined. The garden table costs $98 less than the bench. What is the cost of the bench?

OpenStudy (anonymous):

G+B=652 G=B-98 so replace G from second equation in first (B-98)+B = 652 2B-98 =652 2B=652+98 B=(652+98)/2